OpenStudy (anonymous):

Using the elimination method what are the solutions to this a+8b=8, a+3b=-2

OpenStudy (anonymous):

help me please

OpenStudy (ajprincess):

a+8b=8-(1) a+3b=-2-(2) Subtract equation (2) from equation (1). can u plz tell me vat u get @cdelomas?

OpenStudy (anonymous):

7

OpenStudy (ajprincess):

hmm no. (1)-(2)=> a+8b-a-3b=8-(-2) 5b=8+2 5b=10 Can u find b @cdelomas?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

The answer is 2

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

Divide by 5 both sides to find b.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes its 2

OpenStudy (anonymous):

so there is one solution

OpenStudy (ajprincess):

yup. Nw to find a put 2 in place of b in a+8b=8. can u find a @cdelomas?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you help me find it

OpenStudy (ajprincess):

a+8b=8 a+8(2)=8 a+16=8 Subtract 16 both sides.

OpenStudy (anonymous):

it hard but ill try

OpenStudy (anonymous):

is the answer 2 with one solution

OpenStudy (ajprincess):

hmm no. a is not equal to 2. a+16=8 a=8-16 =-8

OpenStudy (ajprincess):

Getting it?

OpenStudy (anonymous):

so the answer is -8 with one solution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is the solution set

OpenStudy (anonymous):

anas2000 is that the solution -8

OpenStudy (ajprincess):

hmm no. a=-8 and b=2

OpenStudy (anonymous):

yes -8 and 2 is the answer

OpenStudy (anonymous):

so it would be (-8,2)

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

is this a multiple choice question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then yes its -8,2